(1 II. Find eigen values and eigen vectors of A=0 LO 6. 0 2 -3 01 3 2)
Answer #1

I find eigen values to sepaonuofla ? O O AE 0 2 3 6 - 3 solm? then charecterstic et eigenvalue of a equ | A-NII 0 -0 => 2-d 0ill eigen vector corresponding to d = 1 (A-I) = 0 0 x, 0 -> 3 x2 X3 -3 So we have a homogeneous it ly system Gaussian of lineSo 2 ER 23 11 2, x, => 22 o 11 x, => X = x2 co 23 if 13 x - 1 then eigenvector v = iii eigznector corresponding to dz = 2-3iBy Gaussian Elimination - + 0 0 o R, → R,/[-+31) / 3i 3 O 31 0 0 3 0 g o -3 30 o -3 31 R. R2/31 0 0 - R3 R3+3R2 0 1 -( o -3 s(ii) eigencector corresponding to dz=2+31 Similarly part il į - replace then eigen vector is OT 3 =